OpenStudy (anonymous):

An air hockey table has a perimeter of 26 feet. Its area is 36 square feet. What are the dimensions of the air hockey table? ? feet by ? feet

OpenStudy (anonymous):

please explain

OpenStudy (king):

9 feet by 4 feet

OpenStudy (king):

length=9 breadth=4

OpenStudy (anonymous):

ok thanks king

OpenStudy (anonymous):



OpenStudy (king):

ure welcome!!!!

OpenStudy (slaaibak):

2(x + y) = 26 ----> Perimeter. simplified: x+y= 13 x=13 - y Area: x*y = 36 Substituting x: (13 - y) * y = 36 -y^2 + 13y = 36 y^2 - 13y + 36 = 0 (y-9)(y-4) = 0 Therefore y = 4 or y=9 x = 13 - y therefore x = 9 or x equals 4 So x=4 and y=9 or x=9 and y=4